OpenStudy (anonymous):

Suppose that a newly discovered element called centium has three isotopes...? that occur in nature. These are centium-200, centium-203, and centium-209. Assume that these isotopes occur equal in nature. What will be the average atomic mass of this element?

OpenStudy (anonymous):

To find the atomic mass of a mixture of isotopes, you take the weighted average of the three. It is the weighted mean of the three isotopes. (1/3)(200)+(1/3)(203)+(1/3)(209)= weighted average

OpenStudy (trantom):

204. This is because to get an average you add the terms and divide by the number of terms. 200+203+209=612 612/3=204